OpenStudy (anonymous):

12x^2-14=0 solve for x

OpenStudy (anonymous):

online homework is saying that the answer is square root of 42/6

OpenStudy (anonymous):

any sugestions guys?

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

so

OpenStudy (anonymous):

??

OpenStudy (anonymous):

thats the answer, your book might be wrong unless I made a bloo. But I am pretty confident that is correct

OpenStudy (anonymous):

are you sure your question is correct @jammal_a

OpenStudy (anonymous):

hang on this is better solution. 12X^2=14 x^2= 14/12 x=sqrt(14/12)

OpenStudy (anonymous):

but that still gives you not the answer your book presents.

OpenStudy (anonymous):

i will screenshot it.

OpenStudy (anonymous):

@dominator140 @zaisha

OpenStudy (anonymous):

I am not to sure what I am looking at in there mate, I dont even see the equation you presented us with.

OpenStudy (lgbasallote):

if you're still interested in how it was solved... i can show it to you \[12x^2 - 14 = 0\] add 14 to both sides \[12x^2 = 14\] divide both sides by 12 \[x^2 = \frac{14}{12}\] simplify \[x^2 = \frac 76\] take the square root \[x = \sqrt{\frac 76}\] rationalize \[x = \frac{\sqrt 7}{\sqrt 6} \times \frac{\sqrt 6}{\sqrt 6}\] simplify \[x = \frac{\sqrt {42}}6\] did you get that?

OpenStudy (lgbasallote):

tell me if there's something you don't get in the solution @jammal_a